RC Meeting, Tuesday, November 2 , 2021
RC Meeting Information
RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920 There is a 10:00 RC meeting today, 11/02/2021!
- Beam delivery to Halls B,C,D
Hall C Status
Continue production at 42 muA.
- Good running over the last 24 hours. Dummy run on 11/01/2021 swing shift.
Number of good pions is 8400+ (needs update!).
Current Run Plan
- Maintenance day shifted to 11/03/2021. Acceleration retuning firmly scheduled for 11/05/2021-11/07/2021.
- Maintenance day update: beam will be off around 06:00.
- As we need to go to "Restricted Access" the survey will take a little longer. I was assured it will be completed by 10:00 am so people that need access to the Hall should plan for that time tomorrow morning.
- Accelerator swing shift on 11/03/2021 is listed as "Recovery".
- Continue data taking until early Nov
- Maintenance day 11/03/2021, retuning: 11/05-11/07/2021
- Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
- HMS Q3 insulating vacuum spikes generating annoyance alarms that take minutes to recover. https://logbooks.jlab.org/entry/3930175 Noise levels on transducers look OK, but Junaid's plot indicates Q3 is burping: https://logbooks.jlab.org/entry/3932992 .
- Serious HB LHe low liquid level alarms. https://logbooks.jlab.org/entry/3933051 Always seems to start with a subtle, simultaneous increase in level in Q1/2/3. https://logbooks.jlab.org/entry/3929152 .
- Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
- Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
- HCCAM02 not functioning (lift required to replace).
- How the harp scans and carbon hole target should look (see links):
- carbon hole target beam centering procedure:
- Resetting the "ioc heartbeat epics communications" (start by clicking on the heart at upper left of the magnet controls screen
- Fall 2021 Shift Schedule
- Hall-C Live Status
- Hall-C Logbook
- MCC White Board